# Exploring the Deceptive Power of LLM-Generated Fake News: A Study of Real-World Detection Challenges

### v1

@@ 3.2 Human Study
Previous research focused solely on the accuracy of human classification, neglecting instances where individuals lacked confidence in their responses (Jiang et al., 2023; Su et al., 2023). It is vital to explore the cognitive processes behind human decision-making, not just a simple conclusion. To address this, we devised six human study metrics aimed at comprehensively evaluating how specific characteristics of the articles influence the decision-making process of human subjects.
@@
4 Experiment and Analysis
We examine the costs and effectiveness of various article-generation methods in creating deceptive content. Additionally, we investigate patterns in fake news produced by LLMs to aid in the identification of LLM-generated articles by humans.
4.1 Dataset Composition
Essentially, the dataset is composed of three parts: the real news articles, the human-generated fake news articles, and the LLM-generated fake news articles.
The authentic real news articles are collected from authority medical news websites, including ”NIH” (NIH, 2023) and ”WebMD” (WebMD, 2023). The human-generated fake news articles are collected from archives of authority fact-checking websites including ”AFPFactCheck” (AFPFactCheck, 2023), ”CheckYourFact” (CheckYourFact, 2023), ”FactCheck” (FactCheck, 2023), ”HealthFeedback” (HealthFeedback, 2023), ”LeadStories” (LeadStories, 2023), and ”PolitiFact” (PolitiFact, 2023). The publish dates of the articles span from Jan-01-2017 to May-01-2023. The LLM-generated fake news articles are categorized into eight distinct groups, differentiated by their prompt strategies. For the generation of each group, we randomly sampled real news articles to serve as the sources for fake news article generation. The process continued until we reached the target number of fake news articles or until all available real news articles had been utilized. We then reviewed the generated articles. Any generated articles that merely repeated their source articles were considered unqualified as fake news and subsequently excluded. The ratio of qualified-to-generated articles is detailed in Table 1. Following both automatic and manual evaluations, we included approximately 180 generated articles in each category, constrained by the limits of the qualified-to-generated ratio.
4.2 Generation Cost Statistics
Models Count
Success
Rate $\uparrow$
Avg.
Request $\downarrow$
GPT 3.5 180 0.472 4.237
GPT 4 180 0.527 3.795
VLPrompt (different LLMs) Vicuna 160 0.118 16.95
QA 171 0.126 15.87
QA_s 114 0.084 23.81
Baseline Prompts SUMMARY 180 0.580 3.448
ab_role 180 0.647 3.09
Ablation Prompts ab_sem 180 0.478 4.184
real news 1360
Human Written fake news 469
Table 1: Numbers and the costs of articles in the dataset. $\uparrow$ means higher is better; $\downarrow$ the reverse.
Table 1 exhibits basic statistical information about the dataset, including the number of articles and the associated costs for those generated by LLMs. Specifically, the “Success Rate” metric reflects the proportion of successful article generations out of the total attempts. “Average Request” denotes the average number of LLM requests required per successfully generated article. Ideally, if every generated article meets the quality criteria, the average number of requests would be 2 — one for the initial generation and another for the subsequent qualification.
The “Success Rate” is computed from the ratio between the number of generated articles and the used source articles. For instance, Vicuna generated 160 articles out of the 1360 real news articles, resulting in a success rate (the second column) of $160/1360=0.118$. Generating each article requires two requests: one for generation and another for qualification. Consequently, the average number of requests (the third column) needed by Vicuna to produce one qualified article is $2/0.118=16.95$.
Among the generation methods, “Vicuna”, “QA”, and “QA_s” were unable to produce the target of 180 articles from the pool of 1360 real news articles. The low success rate of Vicuna is caused by the contradiction between the scale of the model and the long, complex prompt. A significant portion of the articles generated using “QA” and “QA_s” were discarded in the qualification module. Specifically, “QA_s” is even more costly than “QA”. The reason behind this is that, when LLMs handle context inconsistency issues, they tend to change the modified detail back if the modification is subtle. On the other hand, “ab_role”, “SUMMARY”, and “GPT4” achieved high success rate. Intuitively, modification of the theme leads to larger differences than modification of the details and thus can easily meet the requirement. Besides, the role-play intention injection module seems to lead to a drop in success rates.
4.3 Human Evaluation
Human qualification. We selected 80 articles at random from our dataset and another 80 that the qualification module had deemed unqualified. Each batch of 80 articles included ten articles generated by each of eight different methods. Two human evaluators were assigned the task of comparing these fake news articles with their original, authentic counterparts to determine their similarity. Due to time constraints and limited human resources, it was impractical to manually review all generated articles. Nonetheless, the findings suggest that the qualification module is generally reliable. In the assessment, all 80 articles from the qualified batch and 57 from the unqualified batch were considered correctly categorized. The evaluators found the remaining 23 articles from the unqualified batch confusing, noting that although the differences from the source articles were minor, these distinctions rendered the articles “may be qualified for differences but unqualified as disinformation.” Consequently, this human evaluation process substantiates the reliability of the qualification module.
Model VLPrompt Baseline Prompts Ablations Prompts
GPT 3.5 GPT 4 Vicuna QA QA_s SUMMARY ab_role ab_sem
Correctness $\downarrow$ 0.250 0.436 0.344 0.383 0.313 0.281 0.125 0.563
Intention $\uparrow$ 0.594 0.750 0.531 0.625 0.563 0.656 0.781 0.656
Detail 0.750 0.719 0.469 0.677 0.438 0.844 0.844 0.906
Neutral $\uparrow$ 0.906 0.936 0.906 0.938 1.000 0.875 0.875 0.969
Informative $\uparrow$ 0.706 0.875 0.875 0.931 0.913 0.688 0.813 0.688
Consistent $\uparrow$ 0.781 0.688 0.875 0.844 0.781 0.938 0.938 0.781
Table 2: Human study scores of fake news articles. Rows list metrics used for human study evaluation while the columns indicate news from different sources. The numbers are the average scores of $10$ articles×$\times 2$ people. Bolded and underline results indicate the bests and second bests.
@@ 4.3 Human Evaluation
Deceptive power over human. To assess the quality of the generated articles, human evaluators were initially tasked with determining the authenticity of 90 articles (80 fake news articles and 10 real news articles), discerning whether they were real or fake. Then, they were instructed to compare the 80 fake news articles against their corresponding source articles to evaluate similarities and differences.
@@
The correctness in Table 2 refers to the proportion of fake news articles that were accurately identified by human evaluators. The results in “Correctness” show that “ab_role” was the most effective in misleading participants. Besides, “ab_role” stood out in both “Intention” and “Consistent”. However, this is also a sign that the role-play module does not contribute to the consistency of the contexts. “Detail” indicates that “vicuna” and “QA_s” are inclined to alter article details, potentially leaving the overall theme unchanged. The “Detail” metric for “QA” reveals that, in the absence of guidance, LLMs are prone to posing questions to specific sections of an article. Besides, all the generated articles achieved high scores in “Neutral” and “Informative”, meaning that the LLMs generally retain the original articles’ writing style and specific details, even without explicit instructions to do so.
4.4 Machine Fake News Detection
@@
Detection model introduction. During the experiment, we considered four types of baseline models, fine-tuned PLMs, SOTA fake news detection models, LoRa fine-tuned LLMs, and ChatGPT 3.5. The characteristics of the chosen models are listed below. The fine-tuned models are extracted from the Hugging Face repositories. The SOTA fake news detection model codes were downloaded from corresponding GitHub repositories.
@@
  • Vicuna (Chiang et al., 2023): An open-source LLM fine-tuned mainly with imitation learning from ChatGPT 4. It has been proven to exhibit competitive performance compared to enterprise-level LLMs with relatively small parameter volumes.
  • ChatGPT 3.5 (OpenAI, 2023): The most widely used iteration of OpenAI’s powerful language model. ChatGPT 3.5 turbo API was employed in this research.
Detection model training. We fine-tuned three PLMs, fully trained two SOTA fake news detection models, and fined-tuned two LLMs using LoRa (Hu et al., 2021) and a one-shot prompt strategy. For PLM fine-tuning, we employed a trainable two-layer feed-forward neural network to convert the latent representation of news articles into binary classification results. However, this approach is not suitable for LLMs. Since most LLMs are primarily designed and pre-trained for text generation, it is challenging to regularize their outputs as binary classification results.
To fill this gap, we provided a prompt with an exemplar article-label pair to guide the LLMs during the supervised generation fine-tuning process. This design follows the recommendation of  (Gao et al., 2020) for prompt-based fine-tuning. Within the prompt, the template regularizes the generation format and enables the loss computed only over the conclusion (i.e., “real” or “fake”) drawn by LLM. The example provides supplementary context to assist the LLM in identifying facts/counterfacts in the other articles. This design reflects the real-world scenario in which fact-checkers try to identify if an incoming article is real or fake – the fact-checkers may have access to labels for their archived articles, but may not necessarily be aware of their relations to the incoming article. After the supervised fine-tuning is finished, the model should be able to classify an incoming article by leveraging both the facts in the training data and the relation between the example article and the article to be classified. The detailed prompts are provided in Appendix A.1.
The fine-tuning experiments for PLMs were conducted on a personal computer with 16GB RAM, an i7-11700F/2.50GHz 8 cores CPU, and a GeForce RTX 3090 GPU. We used a batch size of 4, a learning rate of $52E-$ 5, 300 training epochs, and the Adam optimizer. For fine-tuning 7b LLM models with LoRa, we employed a server with 2048GB RAM, AMD EPYC 7742 64-Core CPU, and a Matrox G200eH3 GPU. The batch size was 32, the learning rate was $42E-$ 4, the models were trained for 4 epochs, and the optimizer was AdamW, following the guidelines provided by Hugging Face. The maximum token length for LLMs is set to 1024 due to the limitations in VRAM.
[…]
@@
Comm. LLM ChatGPT 3.5 0.592 0.536 0.770 0.411
Table 3: Fake news classification results of fine-tuned PLMs, SOTA fake news detection models, LoRa fine-tuned LLMs, and chatGPT 3.5 turbo API. In the header, “ACC” means accuracy, “PRC” means precision, and “RCL” means recall. The best is bolded.
Deceptive power over machines. To assess the machine’s power in detecting LLM-generated fake news, we considered three pre-trained language models (PLMs), two 7b LLMs with LoRa (Hu et al., 2021), two state-of-the-art (SOTA) fake news detection models (Zellers et al., 2019; Zhang et al., 2021), and an enterprise-level LLM. The findings indicate that fine-tuned PLMs often misclassify fake news articles as real. Particularly when the PLM, such as FnBERT, has been pre-trained with human-authored fake news. Additionally, DualEmo (Zhang et al., 2021) achieves the highest accuracy, while the fine-tuned Llama2-7b excels in the F1-score. Across the board, all models demonstrate performance metrics below 0.86 in both accuracy and F1 score. However, there is an improvement in the performance of all detection models when trained without human-written fake news, as detailed in Appendix B.1. This suggests the potential benefit of training various models concurrently for the detection of different types of fake news articles.
@@ 4.5 Pattern Discovery in LLM-Generated Fake News
We also investigated the strategies how LLMs alter real news into fake news. We first performed case study by manually comparing the generated articles with their source articles. Then, we utilized word cloud to visualize distinctions ChatGPT 3.5 identifies between the paired fake and real news articles. The discernible patterns highlighted in this analysis will enable humans to recognize news articles generated by LLMs without relying on machine assistance.
Case Study. During the evaluation, human reviewers observed that LLMs often invert opinions rather than altering specific details, as indicated in the case Figure 3. The generated article altered the statement “drinking more water can reduce the risk of heart failure” to “increase the risk of heart failure” (highlighted in red) while preserving the writing style and content irrelevant to the statement (highlighted in green). This deliberate manipulation makes the fake article confusing to both humans and machines, indicating that people may be able to identify LLM-generated fake news by checking the frequency of negative words.
@@
(a) GPT 3.5
(b) GPT 4
(c) Vicuna
(d) QA
(e) QA_s
(f) SUMMARY
(g) ab_role
(h) ab_sem
Figure 4: Word cloud illustrating the differences between LLM-generated articles and their corresponding authentic news pieces.
Patterns During Modification. During the qualification phase, we tasked ChatGPT 3.5 not just with determining whether generated articles differed from their original news sources, but also with explaining its reasons. If some distinctions consistently recur across the majority of fake-real article pairs, words indicating these distinctions would reveal a pattern in how LLMs modify content. In this case, observing the word frequencies could assist in the identification of LLM-generated articles. To pinpoint the most common differences, we filtered out words irrelevant to the differences, like “article”, “first”, and “second”, and displayed the frequencies of the words through word cloud plots.
From the word clouds, we can see that “mention” and “doe mention” (the tokenized version of “does not mention”) frequently appear in all comments. This indicates that one characteristic of the generated article is missing/modifying details. Another major method that all the prompt strategies use to twist the articles is “raise concern”. Similarly, keywords like “potential question”, “criticize”, and “risk” also point to similar methods. Specifically, keywords like “different theme” and “different perspective” appear more in SUMMARY’s word cloud than the others. “specific” and “detail” usually appear in comments for “QA_s” and “QA”, while in the VLPrompt-related comments both of these two types of keywords are frequent. This aligns with our assumption that SUMMARY modifies articles on the global level, “QA” and “QA_s” tend to modify details, while VLPrompt mixes these two types of modification.
@@ 5 Conclusion
We introduced a novel fake news generation strategy, VLPrompt, and conducted extensive tests to evaluate its effectiveness in deceiving humans and machines. Results show ongoing difficulties in detecting fake news. To improve detection, we’ve released a dataset VLPFN containing real news articles, human-created fake news articles, and LLM-generated fake news articles. The dataset, along with the knowledge acquired from our experiments, will support both humans and machines in identifying LLM-generated fake news articles.
@@ A.1 Prompt Detail
Three generation prompts, one qualification prompt, and one one-shot detection prompt are listed below. During the experiments, we also used alternative versions of the generation prompts for comparison. Specifically, “QA_s” is the same as Prompt 6 below except that it is asked to “Raise a question about a detail in the article”. “ab_role” is the version Prompt 7 without Step 2. “ab_sem” is the version Prompt 7 without the requirements in Step 4. Note that we adjusted the original version to remove the usage of human-made false information for fair comparisons. All the prompts below are filled in a single system message, after which there are user prompts providing articles separately.
@@
B.1 Machine Fake News Detection without Human-Written Fake News
@@
Comm. LLM ChatGPT 3.5 0.769 0.510 0.850 0.364
Table 4: Fake news classification results of fine-tuned PLMs, fake news detection models, LoRa fine-tuned LLMs, and chatGPT 3.5 turbo API. In the header, “ACC” means accuracy, “PRC” means precision, and “RCL” means recall. The best is bolded.

### v2

@@ 4.3 Human Evaluation
Deceptive power over human. Previous research focused solely on the accuracy of human classification, neglecting instances where individuals lacked confidence in their responses (Jiang et al., 2023; Su et al., 2023). It is vital to explore the cognitive processes behind human decision-making, not just a simple conclusion. Thus, we devised six human study metrics aimed at comprehensively evaluating how specific characteristics of the articles influence the decision-making process of human subjects.
@@
To assess the quality of the generated articles, human evaluators were initially tasked with determining the authenticity of 90 articles (80 fake news articles and 10 real news articles), discerning whether they were real or fake. Then, they were instructed to compare the 80 fake news articles against their corresponding source articles to evaluate similarities and differences.
@@
The correctness in Table 2 refers to the proportion of fake news articles that were accurately identified by human evaluators. The results in “Correctness” show that “VLPrompt” was the most effective in misleading participants. Besides, “VLPrompt” stood out in both “Intention” and “Consistent”. However, this is also a sign that the role-play module does not contribute to the consistency of the contexts, even when employing more advanced LLMs such as ChatGPT-4. “Detail” indicates that “Vicuna” and “VLPrompt${}_{+r-s}$” are inclined to alter article details, potentially leaving the overall theme unchanged. The “Detail” metric for “QA” reveals that, in the absence of guidance, LLMs are prone to posing questions to specific sections of an article. Besides, all the generated articles achieved high scores in “Neutral” and “Informative”, meaning that the LLMs generally retain the original articles’ writing style and specific details, even without explicit instructions to do so.
4.4 Automatic Fake News Detection
@@
Detection model introduction. During the experiment, we considered eight detection models including four types of baseline models, fine-tuned PLMs, SOTA fake news detection models, LoRA fine-tuned LLMs, and ChatGPT-3.5. The characteristics of the chosen models are listed below. The fine-tuned models are extracted from the Hugging Face repositories. The SOTA fake news detection model codes were downloaded from corresponding GitHub repositories.
@@
  • Vicuna (Chiang et al., 2023): An open-source LLM fine-tuned mainly with imitation learning from ChatGPT-4. It has been proven to exhibit competitive performance compared to enterprise-level LLMs with relatively small parameter volumes.
  • ChatGPT-3.5 (OpenAI, 2023): The most widely used iteration of OpenAI’s powerful language model. ChatGPT-3.5 turbo API was employed in this research.
Detection model training. We fine-tuned three PLMs, fully trained two SOTA fake news detection models, and fined-tuned two LLMs using LoRA (Hu et al., 2021) and a one-shot prompt strategy. For PLM fine-tuning, we employed a trainable two-layer feed-forward neural network to convert the latent representation of news articles into binary classification results. However, this approach is not suitable for LLMs. Since most LLMs are primarily designed and pre-trained for text generation, it is challenging to restrict their outputs as binary classification results.
To fill this gap, we provided a prompt with an exemplar article-label pair to guide the LLMs during the supervised generation fine-tuning process. This design follows the recommendation of  (Gao et al., 2020) for prompt-based fine-tuning. Within the prompt, the template regularizes the generation format and enables the loss computed only over the conclusion (i.e., “real” or “fake”) drawn by LLM. The example provides supplementary context to assist the LLM in identifying facts/counterfacts in the other articles. This design reflects the real-world scenario in which fact-checkers try to identify if an incoming article is real or fake – the fact-checkers may have access to labels for their archived articles, but may not necessarily be aware of their relations to the incoming article. After the supervised fine-tuning is finished, the model should be able to classify an incoming article by leveraging both the facts in the training data and the relation between the example article and the article to be classified. The detailed prompts are provided in Appendix A.1. The fine-tuning experiments for PLMs were conducted with a batch size of 4, a learning rate of $52E-$ 5, 300 training epochs, and the Adam optimizer. For fine-tuning 7b LLM models with LoRA, we employed a batch size of 32 and a learning rate of $42E-$ 4. The models were trained for 4 epochs, and the optimizer was AdamW, following the guidelines provided by Hugging Face.
Type Model ACCnormal-$\uparrow$ F1normal-$\uparrow$ PRCnormal-$\uparrow$ RCLnormal-$\uparrow$
@@
Comm. LLM ChatGPT-3.5 0.592 0.536 0.770 0.411
Table 3: Fake news classification results of automated detectors. Include fine-tuned PLMs, SOTA fake news detection models, LoRA fine-tuned LLMs, and ChatGPT-3.5 turbo API. In the header, “ACC” means accuracy, “PRC” means precision, and “RCL” means recall. The best is bolded.
Deceptive power over automated detectors. To assess the automated detector’s power in detecting LLM-generated fake news, we considered three pre-trained language models (PLMs), two 7b LLMs with LoRA (Hu et al., 2021), two state-of-the-art (SOTA) fake news detection models (Zellers et al., 2019; Zhang et al., 2021), and an enterprise-level LLM. The findings indicate that fine-tuned PLMs often misclassify fake news articles as real. Particularly when the PLM, such as FnBERT, has been pre-trained with human-authored fake news. Additionally, DualEmo (Zhang et al., 2021) achieves the highest accuracy, while the fine-tuned Llama2-7b excels in the F1-score. Across the board, all models demonstrate performance metrics below 0.86 in both accuracy and F1 score. However, there is an improvement in the performance of all detection models when trained without human-written fake news, as detailed in Appendix B.1. This suggests the potential benefit of training various models concurrently for the detection of different types of fake news articles.
@@ 4.5 Pattern Discovery in LLM-Generated Fake News
We also investigated the strategies for how LLMs alter real news into fake news. We first performed case studies by manually comparing the generated articles with their source articles. Then, we utilized word cloud to visualize distinctions ChatGPT-3.5 identifies between the paired fake and real news articles. The discernible patterns highlighted in this analysis will enable humans to recognize news articles generated by LLMs without relying on machine assistance.
Case study. During the evaluation, human reviewers observed that LLMs often invert opinions rather than altering specific details, as indicated in the case Figure 3. The generated article altered the statement “drinking more water can reduce the risk of heart failure” to “increase the risk of heart failure” (highlighted in red) while preserving the writing style and content irrelevant to the statement (highlighted in green). This deliberate manipulation makes the fake article confusing to both humans and automated detectors, indicating that people may be able to identify LLM-generated fake news by checking the frequency of negative words.
@@
(a) VLPrompt${}_{+r}$
(b) VLPrompt${}_{+r-s}$
(c) VLPrompt
(d) GPT 4
(e) Vicuna
(f) QA
(g) QA_s
(h) SUMMARY
Figure 4: Word cloud of the differences between LLM-generated articles and corresponding real news. The top three common modification strategies are: “doesn’t mention,” “raise concern,” and “different perspective.”
Patterns during modification. During the qualification phase, we asked ChatGPT-3.5 not just to determine whether generated articles differed from their original news sources, but also to explain its reasons. If some distinctions consistently recur across the majority of fake-real article pairs, words indicating these distinctions would reveal a pattern in how LLMs modify content. In this case, observing the word frequencies could assist in the identification of LLM-generated articles. To pinpoint the most common differences, we filtered out words irrelevant to the differences, like “article”, “first”, and “second”, and displayed the frequencies of the words through word cloud plots.
From the word clouds, “mention” and “doe mention” (the tokenized version of “does not mention”) frequently appear in all comments. This indicates that one characteristic of the generated article is dropping/modifying details. Another major method that all the prompts use to twist the articles is “raise concern”. Similarly, keywords like “potential question”, “criticize”, and “risk” also point to similar methods. Specifically, keywords like “different theme” and “different perspective” appear more in SUMMARY’s word cloud than the others. “specific” and “detail” usually appear in comments for “QA_s” and “QA”, while in the VLPrompt-related comments both of these two types of keywords are frequent. This aligns with our assumption that SUMMARY modifies articles on the global level, “QA” and “QA_s” tend to modify details, while VLPrompt mixes these two types of modification.
@@ 5 Conclusion
We introduced a novel fake news generation strategy, VLPrompt, and conducted extensive tests to evaluate its effectiveness in deceiving humans and automated detectors. Results show ongoing difficulties in detecting LLM-generated fake news. To improve detection, we’ve released a dataset VLPFN containing real news articles, human-created fake news articles, and LLM-generated fake news articles (including articles generated by VLPrompt and all the other experimented methods). The dataset, along with the knowledge acquired from our experiments, will support both humans and automated detectors in identifying LLM-generated fake news articles.
@@ A.1 Prompt Detail
Four generation prompts, one qualification prompt, and one one-shot detection prompt are listed below. During the experiments, we also used alternative versions of the generation prompts for comparison. Specifically, “VLPrompt${}_{+r}$” is the version VLPrompt with a role-play module as Step 2 as presented in Prompt 8. “VLPrompt${}_{+r-s}$” is the version VLPrompt${}_{+r}$ without the style alignment requirements in Step 4 (Prompt 8). The role-play module lists several objectives, like “cause social panic,” for the LLMs to choose. The role-play technique, recognized for its high success rate as a jailbreak method (Liu et al., 2023; Shen et al., 2023; Hariri, 2023), helps LLMs determine which details to alter to achieve the set objectives. “Jailbreak” here refers to the prompt techniques designed to bypass the ethical checking of LLMs.
“QA_s” is the same as Prompt 6 below except that it is asked to “Raise a question about a detail in the article”. Note that we adjusted the original version of “SUMMARY” and “QA” to remove the usage of human-made false information for fair comparisons. All the prompts below are filled in a single system message, after which there are user prompts providing articles separately.
@@
Figure 8: The prompt for fake news generation (VLPrompt${}_{+r}$). The role-play module is now Step 2.
@@
B.1 Autometic Fake News Detection without Human-Written Fake News
@@
Comm. LLM ChatGPT-3.5 0.769 0.510 0.850 0.364
Table 4: Fake news classification results of fine-tuned PLMs, fake news detection models, LoRA fine-tuned LLMs, and ChatGPT-3.5 turbo API. In the header, “ACC” means accuracy, “PRC” means precision, and “RCL” means recall. The best is bolded.
B.2 Significance Test
To examine if any of the prompt strategies outperform the others, we performed ANOVA (Girden, 1992) over human and automated detectors’ detection metrics for each prompt and found a significance in distribution means (F-value = 2.35, p-value = 0.04). Consequentially, we performed Turkey’s Honestly Significant Difference (HSD) test over each pair of distributions. Table 5 shows the Tukey’s Honestly Significant Difference (HSD) test result. The attributes used for the test are source-wise baseline performances and human study scores. The results show that the fake news generation methods, “QA”, “VLPrompt”, and “SUMMARY” are of different performance distributions against the other models, i.e., the ‘reject” results are “True”. Among the three methods above, “SUMMARY” does not significantly outperform VLPrompt${}_{+r-s}$ while the others do. Therefore, we can conclude that “QA” and “VLPrompt” outperform the others.
group1 group2 meandiff p-adj lower upper reject
QA SUMMARY 0.0977 0.859 -0.1323 0.3278 False
QA VLPrompt 0.0415 0.9981 -0.1885 0.2716 False
QA VLPrompt${}_{+r-s}$ 0.3155 0.0015 0.0854 0.5455 True
QA VLPrompt${}_{+r}$ 0.3602 0.0002 0.1301 0.5902 True
QA gpt4 0.3318 0.0007 0.1018 0.5619 True
QA vicuna 0.3511 0.0003 0.1211 0.5812 True
SUMMARY VLPrompt -0.0562 0.9899 -0.2863 0.1739 False
SUMMARY VLPrompt${}_{+r-s}$ 0.2178 0.0758 -0.0123 0.4478 False
SUMMARY VLPrompt${}_{+r}$ 0.2625 0.0148 0.0324 0.4925 True
SUMMARY gpt4 0.2341 0.0433 0.0041 0.4642 True
SUMMARY vicuna 0.2534 0.0212 0.0234 0.4835 True
VLPrompt VLPrompt${}_{+r-s}$ 0.274 0.0093 0.0439 0.504 True
VLPrompt VLPrompt${}_{+r}$ 0.3187 0.0013 0.0886 0.5487 True
VLPrompt gpt4 0.2903 0.0046 0.0603 0.5204 True
VLPrompt vicuna 0.3096 0.002 0.0796 0.5397 True
VLPrompt${}_{+r-s}$ VLPrompt${}_{+r}$ 0.0447 0.9971 -0.1853 0.2748 False
VLPrompt${}_{+r-s}$ gpt4 0.0164 1.0 -0.2137 0.2464 False
VLPrompt${}_{+r-s}$ vicuna 0.0357 0.9992 -0.1944 0.2657 False
VLPrompt${}_{+r}$ gpt4 -0.0283 0.9998 -0.2584 0.2017 False
VLPrompt${}_{+r}$ vicuna -0.009 1.0 -0.2391 0.221 False
gpt4 vicuna 0.0193 1.0 -0.2108 0.2494 False
Table 5: Tukey’s HSD test over different fake news sources. The test is produced based on automated detector performances and human study evaluations.